How many 3-digit numbers can be formed from digits 1 – 5 if:

a. If repetition is allowed?

b. If repetition is not allowed?


1
Expert's answer
2022-05-24T09:50:58-0400

a.


"5^3=125"

125 3-digit numbers can be formed from digits 1 – 5 if repetition is allowed.


b.


"P(5,3)=\dfrac{5!}{(5-3)!}=60"

60 3-digit numbers can be formed from digits 1 – 5 if repetition is not allowed.